jhonyy9 (jhonyy9):

prove by math induction ,so n is natural ,greater than 2 , a and b natural ,that n=a+b+1 is always true

OpenStudy (ikram002p):

so you wanna prove if a,b are natural then n should be natural as well ? mmm i dont understand what you are trying to prove :o

OpenStudy (agreene):

does \(N \in \{0\}\) and does \(a \ne b\) ?

OpenStudy (ikram002p):

mm you should say does {0}∈ N also does a≠b mmm i cant see its a condition

OpenStudy (agreene):

indeed, so do you count 0 as a member of the naturals? there isnt a defined use of 0 afaik

OpenStudy (agreene):

I suppose it doesnt really matter since its for \(n>2\) and not \(n\ge2\)...

jhonyy9 (jhonyy9):

you need to prove that for every n will exist an a and b for what this equality will be true

OpenStudy (anonymous):

lets assume it is TRUE for n=k, k=a+b+1, where a and b are natural numbers Now, for n =k+1 k+1=(a+b+1)+1 = (a+1)+b+1 Since a and b are natural numbers , (a+1) and b must also be natural numbers Thus, if it holds TRUE for n=k then it also holds TRUE for n=k+1 Now, when k=3 then 3=a+b+1 3=1+1+1 (lets choose a=1 and b=1 where both a and b are natural numbers) 3=3(TRUE) Hence, the statement is TRUE for n>2

OpenStudy (ikram002p):

mmm

jhonyy9 (jhonyy9):

@ikram002p do you like this proof too or what mean ,,mmm" ?

jhonyy9 (jhonyy9):

ty @sauravshakya !

jhonyy9 (jhonyy9):

so than this is like easy how will be proven by math induction too that m= (p-1)/2 +(q-1)/2 +1 is true m greater than 2 p and q are odd numbers - can prove it the same like this above ? ty for your answer

jhonyy9 (jhonyy9):

@sauravshakya opinion please ty

OpenStudy (anonymous):

Since p and q are odd numbers they can be expressed as p=2a+1 and q=2b+1 where a and b are set of whole numbers, (p-1)/2 = (2a+1-1)/2=a (q-1)/2=(2b+1-1)/2=b

OpenStudy (anonymous):

and this will be similar to the first question

jhonyy9 (jhonyy9):

ty

jhonyy9 (jhonyy9):

@ParthKohli your opinion is the same please ?

Parth (parthkohli):

Yup, the proofs look good to me

jhonyy9 (jhonyy9):

ty good luck

OpenStudy (ikram002p):

well i dont think that induction is good proof for the identity of sum of natural numbers is also natural . its more of abstract algebra than being a proof with induction

OpenStudy (ikram002p):

i would say (weak proof ) sorry , but this is my opinian about it note that induction already use the identity of sum of natural number is natural so itmake no sense

jhonyy9 (jhonyy9):

@satellite73 @mathmale - please your opinion about this ? thank you

jhonyy9 (jhonyy9):

@ikram002p how do you think about proof by reductip ad absurdum method ?

jhonyy9 (jhonyy9):

@ganeshie8 @Hero ... your opinion please ? thank you very much

OpenStudy (ikram002p):

idk what is reductip ad absurdum method , but if you wanna proof that sum ofnatural number is also natural , then u gotta use basic things , like grouping (show natural number is a group ) or using well ordering principle

jhonyy9 (jhonyy9):

how and what proof could better to prove this statement ?